OpenStudy (anonymous):

distance between (-2,12) and (0,0)

OpenStudy (anonymous):

LAME : ) LOL Do Your Work <<<<<<<<

sam (.sam.):

\[\sqrt{(x-x _{1})^2+(y-y_{1})^2}\] \[\sqrt{(-2-0)^{2}+(12-0)^{2}}=2\sqrt{37}\]